# Turing Video-based Cognitive Tests to Handle Entangled Concepts

### v1

@@ 5 Entanglement as a mechanism of contextual updating
The quantum theoretical modelling elaborated in Section 4 allows one to draw some interesting conclusions regarding the appearance of entanglement in the combination of natural concepts and, more important, on the nature of this entanglement. We stress, however, that these conclusions are independent of the domain, conceptual or physical, where entanglement is applied. This means that the results obtained in the present article may also shed new light on the nature of physics entanglement.
@@
Secondly, we have seen that, not only the initial state of the composite entity is entangled, but also all coincidence measurements are entangled, in the quantum representation of The Animal Acts situation. This result is due to the violation of the marginal law of Kolmogorovian probability in the cognitive test which forbids concentrating all the entanglement of the state-measurement situation into the state of the composite entity, as we have seen in Section 2.
@@
Thirdly, in each coincidence measurement, all eigenstates are entangled. This result is due to the fact that, in each coincidence measurement, all possible outcomes correspond to combinations of concepts, e.g., The Bear Snorts is itself a combination of the concepts Bear and Snorts, hence it is reasonable to expect that a non-classical mechanism of meaning attribution, similar to The Animal Acts, occurs.
@@
Fourthly, in each coincidence measurement, some eigenstates, which are the final states at the end of the coincidence measurement when a defined outcome is obtained, exhibit a relatively higher degree of entanglement than others, which can exactly be explained by the fact that entanglement captures meaning attribution, hence higher degrees of entanglement correspond to higher meaning attribution, thus higher judgement probabilities. For example, the eigenstates corresponding to The Horse Whinnies, The Horse Snorts, The Tiger Growls and The Cat Meows are the states that exhibit the highest degree of entanglement in the corresponding coincidence measurement. As observed in Section 4, this can be explained as due to the fact that higher meaning is attributed in the combination process for these items. As a matter of fact, these items score the highest probability of being judged as a good example of the conceptual combination The Animal Acts. By contrast, the eigenstates corresponding to The Bear Whinnies, The Horse Meows, The Tiger Whinnies and The Cat Snorts are the states that exhibit the lowest degree of entanglement, close to product states, in the corresponding coincidence measurement. Again this can be explained as due to the fact that lower meaning is attributed in the combination process for these items. As a matter of fact, these items score the lowest probability of being judged as a good example of the conceptual combination The Animal Acts.
@@
Fifthly, by looking at the CHSH factor in Table 1, we notice that the video-based cognitive test violates the CHSH inequality by an amount that exceeds the value $2\sqrt{2}\approx 2.8284$, i.e. Cirel’son’s bound, which is usually believed to be the theoretical limit to represent in Hilbert space the statistical correlations that are observed in Bell-type tests by pushing all entanglement into the state of the composite entity and considering only product measurements. Again, this is due to the fact that the coincidence measurements are actually entangled, rather than product, measurements. As we have argued in [38], independently of the physical or conceptual domain of reference, if one allows entangled measurements, then a Hilbert space representation is possible also for Bell-type tests which violate Cirel’son’s bound.
@@
The above is even more evident if one considers an entire text produced by human language and its meaning relationship with the words (concepts) composing it. Each time a word (concept) is added to a text, one can speak of an ‘updating of contextuality’, and this updating continues to occur until the end of the text that contains all the words. As we have argued in [38, 42], this mechanism of ‘contextual updating’ to attribute meaning has to be carried by an entangled state, because this is exactly how entangled states are formed in the tensor product of Hilbert spaces. In other words, it is these entangled states that accomplish the contextual updating in the mathematical formalism of quantum theory. The deep structural similarities between physical and conceptual domains, suggest that the mechanism of contextual updating could also explain, better than the ‘spooky action at a distance’ mechanism, how entanglement is produced in physics [42].
@@ Acknowledgements
This work was supported by the project “New Methodologies for Information Access and Retrieval with Applications to the Digital Humanities”, scientist in charge S. Sozzo, financed within the fund “DIUM – Department of Excellence 2023–27” and by the funds that remained at the Vrije Universiteit Brussel at the completion of the “QUARTZ (Quantum Information Access and Retrieval Theory)” project, part of the “Marie Sklodowska-Curie Innovative Training Network 721321” of the “European Unions Horizon 2020” research and innovation program, with Diederik Aerts as principle investigator for the Brussels part of the network.

### v2

@@ 6 Entanglement as a mechanism of contextual updating
The quantum theoretical modelling elaborated in Section 5 allows one to draw some interesting conclusions regarding the appearance of entanglement in the combination of natural concepts and, more important, on the nature of this entanglement. We stress, however, that these conclusions are independent of the domain, cognitive or physical, where entanglement is applied, as we have extensively explained in Sections 2 and 3. This means that the results obtained in the present article may also shed new light on the nature of physics entanglement.
@@
Secondly, we have seen that, not only the initial state of the composite entity is entangled, but also all joint measurements are entangled, in the quantum representation of The Animal Acts situation. This result is due to the violation of the marginal law conditions of Kolmogorovian probability in the cognitive test which forbids concentrating all the entanglement of the state-measurement situation into the state of the composite entity, as we have seen in Section 2, where we have also provided arguments to conclude that no-signaling is at place.
@@
Thirdly, in each joint measurement, all eigenstates are entangled. This result is due to the fact that, in each joint measurement, all possible outcomes correspond to combinations of concepts, e.g., The Bear Snorts is itself a combination of the concepts Bear and Snorts, hence it is reasonable to expect that a non-classical mechanism of meaning attribution, similar to The Animal Acts, occurs.
@@
Fourthly, in each joint measurement, some eigenstates, which are the final states at the end of the joint measurement, exhibit a relatively higher degree of entanglement than others, which can exactly be explained by the fact that entanglement captures meaning attribution, hence higher degrees of entanglement correspond to higher meaning attribution, thus higher judgement probabilities. For example, the eigenstates corresponding to The Horse Whinnies, The Horse Snorts, The Tiger Growls and The Cat Meows are the states that exhibit the highest degree of entanglement in the corresponding joint measurement. As observed in Section 5, this can be explained as due to the fact that higher meaning is attributed in the combination process for these items. As a matter of fact, these items score the highest probability of being judged as a good example of the conceptual combination The Animal Acts. By contrast, the eigenstates corresponding to The Bear Whinnies, The Horse Meows, The Tiger Whinnies and The Cat Snorts are the states that exhibit the lowest degree of entanglement, close to product states, in the corresponding joint measurement. Again this can be explained as due to the fact that lower meaning is attributed in the combination process for these items. As a matter of fact, these items score the lowest probability of being judged as a good example of the conceptual combination The Animal Acts.
@@
Fifthly, by looking at the CHSH factor in Table 1, we notice that the video-based cognitive test violates the CHSH inequality by an amount that exceeds the value $2\sqrt{2}\approx 2.8284$, i.e. Cirel’son’s bound, which is usually believed to be the theoretical limit to represent in Hilbert space the statistical correlations that are observed in Bell-type tests by pushing all entanglement into the state of the composite entity and considering only product measurements. Again, this is due to the fact that the joint measurements are actually entangled, rather than product, measurements. As we have illustrated in Sections 2 and 3, independently of the physical or cognitive domain of reference, if one allows entangled measurements, then a Hilbert space representation is possible also for Bell-type tests which violate Cirel’son’s bound (Aerts et al., 2023b, ).
@@
The above is even more evident if one considers an entire text produced by human language and its meaning relationship with the words (concepts) composing it. Each time a word (concept) is added to a text, one can speak of an ‘updating of contextuality’, and this updating continues to occur until the end of the text that contains all the words. As we have argued in Aerts et al., 2023b ; Aerts et al., 2023a , this mechanism of ‘contextual updating’ to attribute meaning has to be carried by an entangled state, because this is exactly how entangled states are formed in the tensor product of Hilbert spaces. In other words, it is these entangled states that accomplish the contextual updating in the mathematical formalism of quantum theory. The deep structural similarities between physical and cognitive domains, suggest that the mechanism of contextual updating could also explain, better than the ‘spooky action at a distance’ mechanism, how entanglement is produced in physics (Aerts et al., 2023a, ).
@@ Acknowledgements
This work was supported by the project “New Methodologies for Information Access and Retrieval with Applications to the Digital Humanities”, scientist in charge S. Sozzo, financed within the fund “DIUM – Department of Excellence 2023–27” and by the funds that remained at the Vrije Universiteit Brussel at the completion of the “QUARTZ (Quantum Information Access and Retrieval Theory)” project, part of the “European Union Marie Sklodowska-Curie Innovative Training Network 721321”, with Diederik Aerts as principle investigator for the Brussels part of the network.